# From Compression to Expansion: A Layerwise Analysis of In-Context Learning

### v1

@@ 5 Bias-variance Decomposition of Task Vectors
Increasing in-context lengths lead to more compressed representations
We first evaluate how the number of demonstrations affects the geometry of ICL representations using layerwise TDNV. The results in Figure 12 show that increasing the number of demonstrations $K$ consistently reduces TDNV across all layers. This indicates that as more demonstrations are provided, the within-task variance of task vectors decreases while the between-task distance increases. This explains why increasing the number of demonstrations improves performance—it leads to more compressed and more distinct representations in the intermediate layers.
Figure 11: Layerwise TDNV under different number of demonstrations $K$.
Figure 12: PCA of task vector as number of demonstrations $K$ increase.
Bias-Variance Decomposition. In Figure 12, we present a PCA visualization of the most compressed layer for two different tasks that share the same query; see an illustrative example in (1). When no demonstrations are provided ($K=0$), both tasks produce the same vectors that reflect the prior of the pretrained model. As $K$ increases, we observe an intriguing phenomenon: $(i)$ different tasks induce task vectors in distinct directions, yet each task follows a consistent direction; $(ii)$ the variance within each task decreases. Based on this observation, we decompose the task vector $\bm{h}_{i,t}(K)$ (where we highlight the dependence on the number of demonstrations $K$ and omit the superscript $(\ell)$) of each instance into the following components
@@
Unlike the classical bias–variance decomposition—where the mean of multiple models often outperforms individual models due to the ensemble effect—the setting here is more nuanced. The mean task vector $\bm{\mu}(K)$ is averaged not only over different demonstrations but also over different queries. Therefore, it is not immediately clear whether the mean task vector still encodes useful task information—and if it does, whether it does more effectively than individual vectors $\bm{h}_{i}(K)$. To address this question, we compare task vector accuracy [9] using both individual task vectors and the mean task vector, with results shown in Figure 15. Remarkably, we observe that injecting the mean task vector consistently leads to better performance, suggesting that it encodes task information more effectively than individual task vectors. Moreover, the performance of the mean task vector also exhibits an inverted U-shaped curve: it peaks around the most compressed layer and improves as the number of demonstrations increases. Together with the results in Figures 12 and 12, this indicates that—ignoring the practical limitations of context length—ICL representations constructed from infinitely many possible demonstrations also exhibit the Layerwise Compression-Expansion phenomenon, and the corresponding task vector $\bm{\mu}(\infty)$ at the most compressed layer can thus be viewed as an ideal task representation. This suggests a promising direction for a theoretical investigation of the phenomenon in the infinite-sample regime, which we leave for future work.
@@
  • Decrease of bias: The task mean vector $\bm{\mu}_{t}(K)$ progressively shifts from the zero-shot mean vector $\bm{\mu}_{t}(0)$ (encodes the pertaining bias) toward the ideal task vector $\bm{\mu}_{t}(\infty)$ as $K$ increases, capturing more information about the task. In other words, the bias term $\bm{\mu}_{t}(K)-\bm{\mu}_{t}(\infty)$ decreases with increasing $K$, and empirically, as shown in Figure 15, we observe that $\|\bm{\mu}_{t}(K)-\bm{\mu}_{t}(\infty)\|_{2}/\|\bm{\mu}_{t}(0)-\bm{\mu}_{t}(\infty)\|_{2}$ decays roughly at a rate of $\mathcal{O}(1/K)$.
  • Decrease of variance: On the other hand, as shown in Figure 15, the variance term $\|\bm{h}_{i,t}(K)-\bm{\mu}_{t}(K)\|_{2}^{2}$ decays roughly at a rate of $\mathcal{O}(1/K)$, which together with the fact that between-task distance becomes a constant when $K$ is large leads to a decay rate of $\mathcal{O}(1/K)$ for TDNV.
@@
Figure 14: Decrease of bias as $\mathcal{O}(1/K)$.
Figure 15: Decrease of variance as $\mathcal{O}(1/K)$.
Theoretical analysis of bias-variance terms for one-layer attention. To develop a theoretical understanding of how the bias and variance terms of task vectors evolve with the number of demonstrations $K$, we consider a simplified setting. Specifically, we analyze a single-layer attention model, as the attention mechanism plays a central role in extracting task representations from demonstrations. To simplify the presentation, we drop the subscript $t$ as we only focus on one task and assume that each demonstration and the query correspond to a single hidden state, denoted as $\bm{h}_{1},\ldots,\bm{h}_{K}\in\mathbb{R}^{d}$ for the demonstrations and $\bm{h}_{q}\in\mathbb{R}^{d}$ for the query, respectively. To facilitate analysis, we adopt a linear attention mechanism, denoted by $\mathrm{Attn}$, which preserves the normalization property of softmax attention, namely, that the attention weights sum to 1 [45, 46]. Linear attention has been widely adopted in the literature for theoretical analysis of ICL [5, 6, 47, 27].
@@
Suppose that each demonstration $\bm{h}_{i},i=1,\ldots,K$ is i.i.d. randomly generated from a distribution $\mathcal{H}$ on $\mathbb{R}^{d}$. Then the output of the query token, $\bm{h}^{\prime}_{q}(K)=[\mathrm{Attn}\left(\bm{h}_{1},\ldots,\bm{h}_{K},\bm{h}_{q})\right)]_{:,K+1}$, where $[]_{:,K+1}$ means the $(K+1)$-th column, satisfies the following statistical properties as the number of demonstration $K$ increases:
@@
The proof for Theorem 5.1 can be found in Appendix D. While our analysis shares the same simplifications and limitations as prior work on linear self-attention  [5, 6, 47, 27], it offers new insights into the functional role of self-attention in ICL, revealing how attention mechanisms contribute to reducing both the variance and bias of task representations—leading to improved performance as the number of demonstrations increases.
@@ 6 Conclusion
This work provides a comprehensive analysis of the internal dynamics of ICL in LLMs. We uncover a prevalent Layerwise Compression–Expansion phenomenon in ICL representations, shedding light on how task information is compressed and later expanded to generate predictions. We show that it has profound implications for ICL performance and robustness and reveal the role of attention mechanisms. These insights not only deepen our understanding of structured representation learning in LLMs but also offer practical implications for improving interpretability, efficiency, and robustness.
@@ 7 Acknowledgement
The appendix is organized as follows. We first provide detailed descriptions of all tasks in Appendix A. We then present additional experiments verifying the Layerwise Compression-Expansion in Appendix B. Next, we detail the saliency map and visualization for all layers in Appendix C. Next, we provide proofs for Theorem 5.1 in Appendix D. Next, we validate the i.i.d condition is important in Theorem 5.1 using the repitition in Appendix E. Finally, we provide detailed description and illustrations of task vector accuracy and early-exit accuracy in Appendix F.
@@ Appendix A Task Description
The detailed descriptions of the tasks used in our empirical studies can be found in Table 1.
Table 1: Descriptions of the tasks used in our empirical studies.
Task Groups Task Example Description Letter-to-Letter Copy Letter a $\to$ a Output the same letter of the given letter. Next Letter a $\to$ b Output the next letter of the given letter in the alphabet. To Uppercase a $\to$ A Output the corresponding uppercase letter of the given lowercase letter. Prev Letter b $\to$ a Output the previous letter of the given letter in the alphabet. Next 2 Letter a $\to$ c Output the letter that comes two positions after the given letter in the alphabet. List-to-Element List First [a,b,c] $\to$ a Output the first item in the given list. List Last [a,b,c] $\to$ c Output the last item in the given list. List Length [a,b,c] $\to$ 3 Output length of the given list. List First Upper [a,b,c] $\to$ A Get the first item in the given list, then output the corresponding uppercase letter. List Last Upper [a,b,c] $\to$ C Get the last item in the given list, then output the corresponding uppercase letter.
Appendix B Additional Experiments
B.1 Universal of Layerwise Compression-Expansion across different tasks
@@
In Figure 1, we validate the Layerwise Compression-Expansion phenomenon across layers on Letter-to-Letter tasks. To further assess its generality, we evaluate the phenomenon on the following task groups: (i) List-to-Element tasks (list-first, list-last, list-length, list-first-upper, list-last-upper); (ii) A combination of Letter-to-Letter and List-to-Element tasks.
@@
We use the DeepSeek-Coder-7B model under a 15-shot ICL setting. As shown in Figure 16, the TDNV exhibits a U-shaped curve across layers in both settings, and both task vector accuracy and early-exit accuracy follow patterns similar to those observed in Figure 1. Notably, Figure 16(b) presents results for the combined task groups listed in Table 1, further supporting the conclusion that this phenomenon holds broadly across diverse tasks. These findings further confirm the universality of the Layerwise Compression-Expansion.
@@
Figure 16: Layerwise Compression-Expansion phenomenon across different tasks groups: a) List-to-Element, b) a combination of two task groups, the Letter-to-Letter and List-to-Element. TDNV first decreases then increases from shallow to deep layers, splitting the model into compression and expansion stages.
Appendix C Saliency Maps
C.1 Definition
Previous sections have extensively explored the use of TDNV to quantify layerwise information compression during ICL, revealing important statistical properties of model representations. In this subsection, we complement the understanding from a fine-grained token level. In particular, we will use the method of saliency maps [40], specifically elucidating which parts of the input significantly contribute to the model’s decision-making. By highlighting critical token interactions, saliency maps provide intuitive insights into model behavior. Denoting by $I_{\ell}$ saliency map at the $\ell$-th layer, we compute it by [40]

### v2

@@ 5 Bias-variance Decomposition of Task Vectors
Increasing in-context lengths lead to more compressed representations.
We first evaluate how the number of demonstrations affects the geometry of ICL representations using layerwise TDNV. The results in Figure˜11 show that increasing the number of demonstrations $K$ consistently reduces TDNV across all layers. This indicates that as more demonstrations are provided, the within-task variance of task vectors decreases while the between-task distance increases. This explains why increasing the number of demonstrations improves performance—it leads to more compressed and more distinct representations in the intermediate layers.
Bias-Variance decomposition. In Figure˜11, we present a PCA visualization of the most compressed layer for two different tasks that share the same query; see an illustrative example in (1). When no demonstrations are provided ($K=0$), both tasks produce the same vectors that reflect the prior of the pretrained model. As $K$ increases, we observe an intriguing phenomenon: $(i)$ different tasks induce task vectors in distinct directions, yet each task follows a consistent direction; $(ii)$ the variance within each task decreases. Based on this observation, we decompose the task vector $\bm{h}_{i,t}(K)$ (where we highlight the dependence on the number of demonstrations $K$ and omit the superscript $(\ell)$) of each instance into the following components
@@
Unlike the classical bias–variance decomposition—where the mean of multiple models often outperforms individual models due to the ensemble effect—the setting here is more nuanced. The mean task vector $\bm{\mu}(K)$ is averaged not only over different demonstrations but also over different queries. Therefore, it is not immediately clear whether the mean task vector still encodes useful task information—and if it does, whether it does more effectively than individual vectors $\bm{h}_{i}(K)$. To address this question, we compare task vector accuracy [8] using both individual task vectors and the mean task vector, with results shown in Figure˜14. Remarkably, we observe that injecting the mean task vector consistently leads to better performance, suggesting that it encodes task information more effectively than individual task vectors. Moreover, the performance of the mean task vector also exhibits an inverted U-shaped curve: it peaks around the most compressed layer and improves as the number of demonstrations increases. Together with the results in Figures˜11 and 11, this indicates that—ignoring the practical limitations of context length—ICL representations constructed from infinitely many possible demonstrations also exhibit the Layerwise Compression-Expression phenomenon, and the corresponding task vector $\bm{\mu}(\infty)$ at the most compressed layer can thus be viewed as an ideal task representation. This suggests a promising direction for a theoretical investigation of the phenomenon in the infinite-sample regime, which we leave for future work.
@@
  • Decrease of bias: The task mean vector $\bm{\mu}_{t}(K)$ progressively shifts from the zero-shot mean vector $\bm{\mu}_{t}(0)$ (encodes the pertaining bias) toward the ideal task vector $\bm{\mu}_{t}(\infty)$ as $K$ increases, capturing more information about the task. In other words, the bias term $\bm{\mu}_{t}(K)-\bm{\mu}_{t}(\infty)$ decreases with increasing $K$, and empirically, as shown in Figure˜14, we observe that $\|\bm{\mu}_{t}(K)-\bm{\mu}_{t}(\infty)\|_{2}/\|\bm{\mu}_{t}(0)-\bm{\mu}_{t}(\infty)\|_{2}$ decays roughly at a rate of $\mathcal{O}(1/K)$.
  • Decrease of variance: On the other hand, as shown in Figure˜14, the variance term $\|\bm{h}_{i,t}(K)-\bm{\mu}_{t}(K)\|_{2}^{2}$ decays roughly at a rate of $\mathcal{O}(1/K)$, which together with the fact that between-task distance becomes a constant when $K$ is large leads to a decay rate of $\mathcal{O}(1/K)$ for TDNV.
@@
Figure 13: Decrease of bias at rate of $\mathcal{O}(1/K)$.
Figure 14: Decrease of variance at rate of $\mathcal{O}(1/K)$.
Theoretical analysis of bias-variance terms for attention layer. To develop a theoretical understanding of how the bias and variance terms of task vectors evolve with the number of demonstrations $K$, we consider a simplified setting. Specifically, we analyze a single-layer attention model, as the attention mechanism plays a central role in extracting task representations from demonstrations. To simplify the presentation, we drop the subscript $t$ as we only focus on one task and assume that each demonstration and the query correspond to a single hidden state, denoted as $\bm{h}_{1},\ldots,\bm{h}_{K}\in\mathbb{R}^{d}$ for the demonstrations and $\bm{h}_{q}\in\mathbb{R}^{d}$ for the query, respectively. To facilitate analysis, we adopt a linear attention mechanism, denoted by $\mathrm{Attn}$, which preserves the normalization property of softmax attention, namely, that the attention weights sum to 1 [46, 47]. Linear attention has been widely adopted in the literature for theoretical analysis of ICL [5, 24, 48, 28].
@@
Suppose that each demonstration $\bm{h}_{i},i=1,\ldots,K$ is i.i.d. randomly generated from a distribution $\mathcal{H}$ on $\mathbb{R}^{d}$. Then the output of the query token, $\bm{h}^{\prime}_{q}(K)=[\mathrm{Attn}\left(\bm{h}_{1},\ldots,\bm{h}_{K},\bm{h}_{q})\right)]_{:,K+1}$, where $_{:,K+1}$ means the $(K+1)$-th column, satisfies the following statistical properties as the number of demonstration $K$ increases:
@@
The proof for Theorem˜5.1 can be found in Appendix˜E. While our analysis shares the same simplifications and limitations as prior work on linear self-attention  [5, 24, 48, 28], it offers new insights into the functional role of attention in ICL, revealing how attention contribute to reducing both the variance and bias of task representations—leading to improved performance as the number of demonstrations increases.
Figure 15: Layerwise TDNV using model trained with CE loss v.s. CE + contrastive loss on layer 77.
Figure 16: Task-vector contrastive fine-tuning improves task-vector accuracy.
6 Applications of Compression-to-Expression
Identify the optimal task vector layer. To find the optimal intermediate layer for task vector extraction, previous works [8] typically patch the vector at each layer and evaluate accuracy using a validation set. TDNV provides a more efficient way to identify the optimal layer for task vectors. As shown in Figure˜1, at a certain intermediate layer $\hat{\ell}$, we observe three simultaneous changes: the TDNV shifts from decreasing to increasing, the task vector accuracy begins to decrease, and the early-exit accuracy starts to increase. Since the layer with minimum TDNV corresponds to the layer with maximum task vector accuracy, we can identify the optimal layer with just one pass of inference using $\hat{\ell}=\operatorname*{\text{arg~min}}_{\ell\in[1,L]}\text{TDNV}^{(\ell)}$.
Task-vector contrastive fine-tuning improves task-vector accuracy. Motivated by prior evidence that more compressed task vectors yield better performance, we propose task-vector contrastive fine-tuning that explicitly encourages such compression. Specifically, during fine-tuning on ICL tasks, we augment the cross-entropy (CE) loss with a contrastive loss applied to the task vectors. This loss pulls representations from the same task closer together while pushing apart those from different tasks (see Appendix˜H for the exact formulation). We fine-tune a pretrained GPT-2 model on symbolic ICL domains using either the baseline CE loss or our combined loss applied in $7$-th layer. As shown in Figure˜16, the contrastive term lowers TDNV, indicating stronger task-vector compression, which in turn boosts downstream task-vector accuracy by an average of 20% (Figure˜16).
@@ 7 Conclusion
This work provides a comprehensive analysis of the internal dynamics of ICL in LLMs. We uncover a prevalent Layerwise Compression-Expression phenomenon in ICL representations, shedding light on how task information is compressed and later expressed to generate predictions. We show that it has profound implications for ICL performance and robustness and reveal the role of attention mechanisms. These insights not only deepen our understanding of structured representation learning in LLMs but also offer practical implications for improving interpretability, efficiency, and robustness.
@@ 8 Acknowledgement
The appendix is organized as follows. We first provide detailed descriptions of all tasks in Appendix˜A. Next, we add more discussion on related works in Appendix˜B. Appendix˜C presents additional experiments, including verification of the Layerwise Compression-Expression and an ablation study on the choice of the last separation tokens. Appendix˜D offers a fine-grained token-level analysis, featuring saliency maps and grid TDVN visualizations. Proofs of Theorem˜5.1 are given in Appendix˜E, followed by an examination of the i.i.d. assumption through repetition experiments in Appendix˜F. We then provide detailed descriptions and illustrations of task-vector accuracy and early-exit accuracy in Appendix˜G. Finally, Appendix˜H presents illustrations of the task-vector contrastive fine-tuning method and PCA visualization of extracted task vectors.
@@ A.1 Symbolic Tasks
Detailed descriptions of the symbolic tasks used in our empirical studies are provided in Table˜1, covering the algorithmic, translation, linguistic, and knowledge domains.
Table 1: Descriptions of symbolic tasks.
Task Domains Task Example Description Algorithmic(Letter-to-Letter) Copy Letter a $\to$ a Output the same letter of the given letter. Next Letter a $\to$ b Output the next letter of the given letter in the alphabet. To Uppercase a $\to$ A Output the corresponding uppercase letter of the given lowercase letter. Prev Letter b $\to$ a Output the previous letter of the given letter in the alphabet. Next 2 Letter a $\to$ c Output the letter that comes two positions after the given letter in the alphabet. Algorithmic(List-to-Element) List First [a,b,c] $\to$ a Output the first item in the given list. List Last [a,b,c] $\to$ c Output the last item in the given list. List Length [a,b,c] $\to$ 3 Output length of the given list. List First Upper [a,b,c] $\to$ A Get the first item in the given list, then output the corresponding uppercase letter. List Last Upper [a,b,c] $\to$ C Get the last item in the given list, then output the corresponding uppercase letter. Translation French $\to$ English bonjour $\to$ hello Translate the given French word into English. Spanish $\to$ English gracias $\to$ thank you Translate the given Spanish word into English. English $\to$ French goodbye $\to$ au revoir Translate the given English word into French. English $\to$ Italian music $\to$ musica Translate the given English word into Italian. English $\to$ Spanish thank you $\to$ gracias Translate the given English word into Spanish. Linguistic Antonyms hot $\to$ cold Output the antonym of the given word. Plural $\to$ Singular cats $\to$ cat Convert the given plural noun to its singular form. Present Simple $\to$ Gerund run $\to$ running Convert the given verb from present simple to its gerund form. Present Simple $\to$ Past Perfect walk $\to$ had walked Convert the given verb from present simple to past perfect tense. Present Simple $\to$ Past Simple jump $\to$ jumped Convert the given verb from present simple to past simple tense. Singular $\to$ Plural dog $\to$ dogs Convert the given singular noun to its plural form. Knowledge Country $\to$ Capital France $\to$ Paris Output the capital city of the given country. Football Player $\to$ Position Lionel Messi $\to$ Forward Output the playing position of the given football player. Location $\to$ Continent Brazil $\to$ South America Output the continent where the given location is found. Location $\to$ Country Kyoto $\to$ Japan Output the country in which the given location is situated. Location $\to$ Language Egypt $\to$ Arabic Output the primary language spoken in the given location. Location $\to$ Religion India $\to$ Hinduism Output the predominant religion of the given location.
A.2 Language Understanding Tasks
For evaluating TDNV on natural language understanding tasks, we require each query sentence to be assessed across multiple attributes. To enable this, we construct a synthetic natural language dataset in which every sentence can be evaluated on six distinct attributes: length, semantic polarity, tense, sentence type, subject person, and entity type. Each attribute is associated with several categorical labels, all of which are summarized in Table˜2. The dataset contains 1,000 samples, and representative examples of these sentences are provided in Table˜3.
Attribute Labels Length short, medium, long Semantic Polarity positive, negative, neutral Tense present, past, future, progressive Sentence Type declarative, interrogative, imperative, exclamatory Subject Person first_person, second_person, third_person Entity Type person, location, organization
Table 2: The attributes and labels in language understanding dataset.
Sentence Length Semantic Polarity Tense Sentence Type Subject Person Entity Type I enjoy morning walks. short positive present declarative first_person person Close the window now. short neutral present imperative second_person location Despite the heavy rain, our research team successfully completed the outdoor experiment and gathered all the required samples before sunset. long positive past declarative third_person organization Will you be visiting the United Nations headquarters in New York next year to attend the global climate summit? long neutral future interrogative second_person location While the orchestra rehearsed the challenging new symphony, the conductor meticulously adjusted each section to achieve the perfect balance of sound for the upcoming performance. long neutral progressive declarative third_person organization
Table 3: Example sentences in language understanding dataset, each sentence is annotated with 6 attributes.
A.3 Multimodality Tasks
For evaluating TDNV on multimodality tasks, we require each query image to be assessed across multiple attributes. To enable this, we construct a synthetic vision-text dataset in which every image can be evaluated on four distinct attributes: color, shape, size and texture. Each attribute is associated with several categorical labels, all of which are summarized in Table˜4. The dataset contains 300 samples, and representative examples of these images are provided in Figure˜17.
Attribute Labels Color red, green, blue, yellow, black Shape circle, square, triangle, pentagon, star Size small, medium, large Texture solid, stripes, dots, checker
Table 4: The attributes and labels in multimodality dataset.
Figure 17: Example images in multimodality dataset, each image is annotated with 4 attributes.
Appendix B Related Works
Layerwise Representations An intriguing line of research [17, 18, 19, 20, 21, 22] has empirically investigated the role of different layers in feature learning. These studies show that in image classification tasks, features in intermediate layers become increasingly linearly separable as the layers deepen. Specifically, Neural Collapse ($\mathcal{NC}$) properties emerge in intermediate layers, where the within-class variance decreases compared to the between-class variance as depth increases. This indicates that layerwise compression occurs monotonically with layer depth in these settings. However, our hypothesis reveals that in the ICL setting, decoder-only models’ layerwise representations exhibit dual encoding-decoding stages: shallow layers compress information while deep layers express it. Furthermore, research by [49] shows that intermediate layers consistently outperform both shallow and final layers on downstream tasks. Our research focuses on how intermediate layers achieve maximum information compression in the ICL setting.
In-Context Learning Interpretability Numerous studies have focused on the working mechanisms of ICL [3, 4, 5, 23, 24, 6]. Xie et al. [3] propose that ICL emerges as an implicit form of Bayesian inference. In the realm of meta-learning, Chen et al. [4] introduce in-context tuning to predict target labels from concatenated sequences of task instructions. A significant line of research connects ICL with gradient descent optimization. Von Oswald et al. [5] demonstrate that Transformers trained on autoregressive tasks can emulate gradient descent in their forward pass. Dai et al. [23], Ahn et al. [24] compare standard gradient descent-based fine-tuning with ICL, revealing that transformer attention in ICL exhibits a dual form of gradient descent-based optimization. Olsson et al. [6] identify "induction heads"—specific attention heads that facilitate ICL by copying patterns from previous tokens. However, our work focuses on the layer-wise representational analysis of ICL. In addition, Doimo et al. [26] also reveals different clustering patterns in shallow and deep layers. However, we observed a strikingly different phenomenon and would like to clarify the differences in settings and results. Doimo et al. [26] uses demonstrations with minimal task cues while the query itself provides ample task information, enabling high zero-shot accuracy and only modest few-shot gains. Our tasks encode the task only in the demonstration pairs, while the query provides no clues, yielding near-random zero-shot performance. As a result, our work shows that the model progressively compresses task information from the demonstration pairs, reaching maximal compression around the intermediate layers. In contrast, Doimo et al. [26] clusters representations by semantic subject, with ARI peaking in the earliest layers.
Appendix C Additional Experiments
C.1 Universal of Layerwise Compression-Expression across different tasks
@@
In Figure˜1, we validate the Layerwise Compression-Expression phenomenon across layers on Letter-to-Letter tasks. To further assess its generality, we evaluate the phenomenon on the other algorithmic task groups: (i) List-to-Element tasks (list-first, list-last, list-length, list-first-upper, list-last-upper); (ii) A combination of Letter-to-Letter and List-to-Element tasks.
@@
We use the DeepSeek-Coder-7B model under a 15-shot ICL setting. As shown in Figure˜18, the TDNV exhibits a U-shaped curve across layers in both settings, and both task vector accuracy and early-exit accuracy follow patterns similar to those observed in Figure˜1. Notably, Figure˜18(b) presents results for the combined task groups listed in Table˜1, further supporting the conclusion that this phenomenon holds broadly across diverse tasks. These findings further confirm the universality of the Layerwise Compression-Expression.
@@
Figure 18: Layerwise Compression-Expression phenomenon across different tasks groups: a) List-to-Element, b) a combination of two task groups, the Letter-to-Letter and List-to-Element. TDNV first decreases then increases from shallow to deep layers, splitting the model into compression and expression stages.
C.2 Alternative Representations for task vector
We choose the last separator token as the ICL representation following prior work [8], where it serves as a natural anchor point between the demonstrations and the query. For comparison, we evaluate two other aggregation strategies: ($i$) Mean of All Tokens: the mean of all demonstration token representations(remove the last query and separator), and ($ii$) Mean of All Separator Tokens: the mean of all separator token representations. To quantitatively evaluate which representation best captures layerwise features, we compare the TDNV changes during both compression and expression stages:
$\Delta_{\text{Compression}}=\frac{\mathrm{TDNV}_{0}-\mathrm{TDNV}_{\ell_{\min}}}{\mathrm{TDNV}_{0}}\qquad\text{and}\qquad\Delta_{\text{Expression}}=\frac{\mathrm{TDNV}_{L}-\mathrm{TDNV}_{\ell_{\min}}}{\mathrm{TDNV}_{L}}.$
Where $\mathrm{TDNV}_{0}$ and $\mathrm{TDNV}_{L}$ are the TDNV of the first and last layer, and $\mathrm{TDNV}_{\ell_{\min}}$ is the minimum TDNV.
Representation Compression Ratio $\uparrow$ Expression Ratio $\uparrow$
Mean All Tokens 0.9671 0.6002
Mean Sep Tokens 0.9879 0.5448
Last Sep Token 0.9926 0.7746
Table 5: Compression and expression ratios for different representations.
As shown Table˜5, the Last Sep Token representation demonstrates both higher compression and expression ratios. This evidence supports our conclusion that the last separator token remains the optimal choice for capturing task-relevant information.
The other two alternatives are suboptimal for the following reasons:
  • Mean of All Tokens. As shown in Figure˜20, saliency maps reveal that token contributions to task representation are highly uneven. Early layers focus on label tokens within the demonstrations, while later layers shift attention to the final separator token as the primary carrier of task information. Averaging across all tokens therefore introduces noise from irrelevant content tokens and dilutes the in-context learning (ICL) signal.
  • Mean of All Separator Tokens. As illustrated in Figure˜21, grid-level TDNV analysis across separator tokens shows a monotonic decrease in TDNV from the first to the last separator. This pattern indicates that the model progressively compresses task-relevant information across successive demonstrations. Because later separators encode richer context and stronger compression, averaging over all separators weakens this effect, whereas using only the final separator captures the fully accumulated task representation.
Appendix D Token Level Analysis
D.1 Saliency Maps
In the main pages, we have extensively explored the use of TDNV to quantify layerwise information compression during ICL, revealing important statistical properties of model representations. In this subsection, we complement the understanding from a fine-grained token level. In particular, we will use the method of saliency maps [50], specifically elucidating which parts of the input significantly contribute to the model’s decision-making. By highlighting critical token interactions, saliency maps provide intuitive insights into model behavior. Denoting by $I_{\ell}$ saliency map at the $\ell$-th layer, we compute it by,